# Diagonalization Procedure for Reducible Massless ${\cal N}=1$ Supermultiplets

### v2

@@ 4 Diagonalization for $n=2$
The diagonalization procedure for the case $n=2$ essentially repeats the same steps as for $n=1$. However it has some novel features, such as a presence of the simplest mixed symmetry fields.
@@
As it was in the case of $n=1$, the precise from of this decomposition is obtained from the requirement that the Lagrangian splits into a sum of Lagrangians for individual Fronsdal modes. However one can get the same result by requiring that the fields $\Phi_{\mu_{1}\mu_{2}\mu_{3}}$ and $\Phi_{\mu}$ have “proper” gauge transformation rules i.e.,
@@
For system describing mixed symmetry fields one perform the following field redefinitions
@@
This transformation results into a decomposition of the Lagrangian (4.8) into a sum of a standard Lagrangian for a massless vector field $T_{\mu}$
@@
This Lagrangian can be brought into a sum of Lagrangians, each describing an individual Fang-Fronsdal mode by using the transformations
@@
Let us make a few comments on $AdS_{d}$ deformation for the case of $n=2$. Recall, that according to the so-called BMV conjecture [43], performing an $AdS_{d}$ deformation of the Lagrangian (4.17) one loses the gauge invariance with respect to the symmetric parameter $\Lambda_{\mu\nu}$ and as the result the number of on-shell degrees of freedom changes [^5] See [44][46] for a proof of the BMV conjecture, [47] for a discussion for the case of the de Sitter background and [48] for recent progress in the BRST approach.. In order to avoid this issue, one introduces an extra Stueckelberg-type rank-two field, which couples to the original mixed symmetry field in the Lagrangian. This new field transforms under the parameter $\Lambda_{\mu\nu}$ and also brings “ its own” gauge symmetry with a parameter $\xi_{\mu}$. Then one can prove, that after fixing all gauge symmetries, the modified system contains the same degrees of freedom as the original Lagrangian on the flat background. From the consideration above one can conclude, that in order to repeat the procedure described in Subsection 3.2 for the case at hand, one has to follow the option when one performs the splitting of the original Lagrangian into irreducible representations prior to the $AdS_{d}$ deformation. In this case a coupling of the Stueckelberg-like field to the symmetric part of $\phi_{\mu_{1}\mu_{2},\nu}$ can be avoided, and this keeps a possibility for the diagonalization procedure also on $AdS_{d}$. A detailed study of this question, is beyond the scope of the present paper and we leave it for a future work.
@@ 5 Diagonalization for a general $n$
The procedure that was shown in detail in the previous two Sections can be generalized to the case of an arbitrary $n$. Using the language of auxiliary oscillators, described in the Section 2, one starts with the Lagrangian (2) and splits the field $|\phi\rangle$ into a sum
@@
where the vectors $|\tilde{D}\rangle$ and $|\tilde{M}\rangle$ have the form similar to (5.2), namely
@@
Inverting these transformations
@@
one can check that in terms of the newly introduced fields the Lagrangian has a form (3.3), where
@@
The expansion coefficients $c_{q}$ can be found from the requirement of the double traselesness condition on $|\Phi_{n}\rangle$, which should be satisfied by a Fronsdal mode. Alternatively, most of the coefficients $c_{q}$ (except the ones that multiply gauge-invariant combinations $|{\cal W}_{n}^{0}\rangle$) can also be found from the requirement that the Fronsdal field has the “proper” gauge transformations $\delta|\Phi_{n}\rangle=l^{+}|\Lambda_{n-1}\rangle$, with the traceless gauge parameter. In order to check, that the Lagrangian for reducible higher spin fields splits into a sum of the Lagrangians, each describing an irreducible mode, one has to invert the transformations (5.11) and insert the inverted expression into the Lagrangian (5.9). The fact, that this procedure diagonalizes the Lagrangian was explicitly checked up to the spin four, but as it follows from the discussion above, it is natural to expect, that the diagonalization of Lagrangians works for a general value of $n$.
@@
one looks for a solution for an irreducible mode $|\tilde{\Psi}_{n}\rangle$ in the form of expansion
@@
For mixed symmetry fields, the procedure is essentially the same, with the exception that now one has two types of ${\cal W}$, due to the two possible types symmetries contained in the field $|\tilde{M}\rangle$. Bellow we shall illustrate this procedure for the case when the field $|{\cal B}\rangle$ is of the type $Y(3,1)$, a generalization to an arbitrary $n$ is obvious.
@@
The parameter $\Xi_{\mu_{1}\mu_{2},\nu}$ is symmmetrical with respect to its first two indices, but has no symmetry with respect to the third index. Notice also that the gauge transformation rule (5.16) is consistent with the symmetry of the field ${\cal B}_{\mu_{1}\mu_{2}\mu_{3},\nu}$ i.e., with the condition $t|{\cal B}\rangle=0$. Next we define in analogy with (5.10)
@@ 6 Conclusion
There are several directions in which our results may be applied. In particular, the diagonalization procedure seems to be essential tool in the construction of supersymmetric Lagrangians which contain reducible higher spin fields on an $AdS_{d}$ background, a problem that we hope to address in a separate publication. It would also be interesting to apply the diagonalization procedure developed in the present paper to supersymmetric models which contain cubic and higher-order interactions [49][52] either on flat or on $AdS_{d}$ backgrounds similar to how it has been done for the case of current-current interactions [27][28].
@@ Appendix A Lagrangians and the algebra of operators on AdS space
Although in this paper we were primarily concerned with a flat background, here we give the description of massless bosonic and fermionic higher spin fields for a $d$-dimensional anti-de Sitter space, originally given for $d=4$ in [56][57]. The corresponding equations for the Minkowski space-time can be simply recovered by taking the limit $L\rightarrow\infty$, where $L$ is the radius of the $AdS_{d}$ space.
@@
$t|\tilde{\Phi}\rangle\rightarrow\tilde{\Phi}_{(\mu_{1},...,\mu_{s},\nu)}(x),\qquad\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,t^{+}|\Phi\rangle\rightarrow\Phi_{\mu_{1},...,\nu}(x)$ (A.29)
$M_{12}|\tilde{\Phi}\rangle\rightarrow\tilde{\Phi}^{\nu}{}_{\mu_{2},...,\mu_{s},\nu}(x),\qquad\,\,\,\,\,\,M_{12}^{+}|\Phi\rangle\rightarrow g_{\nu(\mu_{1}}\Phi_{\mu_{2},...,\mu_{s})}(x)$ (A.30)

### v3

@@ 4 Diagonalization for $n=2$
The diagonalization procedure for the case $n=2$ essentially repeats the same steps as for $n=1$. However, it has some novel features, such as presence of the simplest mixed symmetry fields.
@@
As it was in the case of $n=1$, the precise from of this decomposition is obtained from the requirement that the Lagrangian splits into a sum of Lagrangians for individual Fronsdal modes. However, one can get the same result by requiring that the fields $\Phi_{\mu_{1}\mu_{2}\mu_{3}}$ and $\Phi_{\mu}$ have “proper” gauge transformation rules i.e.,
@@
For the system describing mixed symmetry fields one performs the following field redefinitions
@@
This transformation results in decomposition of the Lagrangian (4.8) into a sum of a standard Lagrangian for a massless vector field $T_{\mu}$
@@
This Lagrangian can be converted into a sum of Lagrangians, each describing an individual Fang-Fronsdal mode by using the transformations
@@
Let us make a few comments on $AdS_{d}$ deformation for the case of $n=2$. Recall, that according to the so-called BMV conjecture [45], performing an $AdS_{d}$ deformation of the Lagrangian (4.17) one loses the gauge invariance with respect to the symmetric parameter $\Lambda_{\mu\nu}$ and as a result the number of on-shell degrees of freedom changes [^6] See [46][48] for a proof of the BMV conjecture, [49] for a discussion for the case of the de Sitter background and [50] for recent progress in the BRST approach.. In order to avoid this issue, one introduces an extra Stueckelberg-type rank-two field, which couples to the original mixed symmetry field in the Lagrangian. This new field transforms under the parameter $\Lambda_{\mu\nu}$ and also brings “ its own” gauge symmetry with a parameter $\xi_{\mu}$. Then one can prove, that after fixing all gauge symmetries, the modified system contains the same degrees of freedom as the original Lagrangian on the flat background. From the consideration above one can conclude, that in order to repeat the procedure described in Subsection 3.2 for the case at hand, one has to follow the option when one performs the splitting of the original Lagrangian into irreducible representations prior to the $AdS_{d}$ deformation. In this case a coupling of the Stueckelberg-like field to the symmetric part of $\phi_{\mu_{1}\mu_{2},\nu}$ can be avoided, and this keeps a possibility for the diagonalization procedure also on $AdS_{d}$. A detailed study of this question, is beyond the scope of the present paper and we leave it for a future work.
@@ 5 Diagonalization for a general $n$
The procedure that was shown in detail in the previous two Sections can be generalized to the case of an arbitrary $n$. Using the language of auxiliary oscillators, described in Section 2, one starts with the Lagrangian (2) and splits the field $|\phi\rangle$ into a sum
@@
where the vectors $|\tilde{D}\rangle$ and $|\tilde{M}\rangle$ have a form similar to (5.2), namely
@@
After inverting these transformations
@@
one can check that, in terms of the newly introduced fields, the Lagrangian has a form (3.3), where
@@
The expansion coefficients $c_{q}$ can be found
$c_{p}=(-1)^{p}\frac{(d+2(s-p-3))!!}{(d+2(s-3))!!}$ (5.12)
from the requirement of the zero double trace condition on $|\Phi_{n}\rangle$, which should be satisfied by a Fronsdal mode. Alternatively, most of the coefficients $c_{q}$ (except the ones that multiply gauge-invariant combinations $|{\cal W}_{n}^{0}\rangle$) can also be found from the requirement that the Fronsdal field has the “proper” gauge transformations $\delta|\Phi_{n}\rangle=l^{+}|\Lambda_{n-1}\rangle$, with the traceless gauge parameter. In order to check that the Lagrangian for reducible higher spin fields splits into a sum of the Lagrangians, each describing an irreducible mode, one has to invert the transformations (5.11) and insert the inverted expression into the Lagrangian (5.9). The fact that this procedure diagonalizes the Lagrangian was explicitly checked up to spin four. However, as it follows from the discussion above, it is natural to expect that the diagonalization of Lagrangians works for a general value of $n$.
@@
one looks for a solution for an irreducible mode $|\tilde{\Psi}_{n}\rangle$ in the form of the expansion
@@
For mixed symmetry fields, the procedure is essentially the same, with the exception that now one has two types of ${\cal W}$, due to the two possible types symmetries contained in the field $|\tilde{M}\rangle$. Bellow we shall illustrate this procedure for the case when the field $|{\cal B}\rangle$ is of the type $Y(3,1)$, a generalization to an arbitrary $n$ is obvious. Recall, that as we discussed above, in order to find the coefficients $c_{p}$ for the case of totally symmetric bosonic and fermionic fields, one can use the conditions of zero double trace and zero triple gamma–trace, respectively. The result is given by equations (5.12) and (5.16) by taking $n=3$. On the other hand, for the mixed symmetry field ${\cal B}_{\mu_{1}\mu_{2}\mu_{3},\nu}$ the zero double trace condition is automatically satisfied and therefore we proceed with the requirement of the “proper” gauge transformations for it.
@@
The parameter $\Xi_{\mu_{1}\mu_{2},\nu}$ is symmetrical with respect to its first two indices, but has no symmetry with respect to the third index. Notice also that the gauge transformation rule (5.17) is consistent with the symmetry of the field ${\cal B}_{\mu_{1}\mu_{2}\mu_{3},\nu}$ i.e., with the condition $t|{\cal B}\rangle=0$. Next we define in analogy with (5.10)
@@ 6 Conclusion
There are several directions in which our results may be applied. In particular, the diagonalization procedure seems to be an essential tool in the construction of supersymmetric Lagrangians which contain reducible higher spin fields on an $AdS_{d}$ background, a problem that we hope to address in a separate publication. It would also be interesting to apply the diagonalization procedure developed in the present paper to supersymmetric models which contain cubic and higher-order interactions [51][54] either on flat or on $AdS_{d}$ backgrounds similar to how it has been done for the case of current-current interactions [27][28].
@@ Appendix A Lagrangians and the algebra of operators on AdS space
Although in this paper we were primarily concerned with a flat background, here we give the description of massless bosonic and fermionic higher spin fields for a $d$-dimensional anti-de Sitter space, originally given for $d=4$ in [58][59]. The field equations for general mixed symmetry fields on $AdS_{d}$ are given in [60]. The corresponding equations for the Minkowski space-time can be simply recovered by taking the limit $L\rightarrow\infty$, where $L$ is the radius of the $AdS_{d}$ space.
@@
$t|\tilde{\Phi}\rangle\rightarrow\tilde{\Phi}_{(\mu_{1},...,\mu_{s},\nu)}(x),\qquad\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,t^{+}|\Phi\rangle\rightarrow\Phi_{\mu_{1},...,\nu}(x)$
$M_{12}|\tilde{\Phi}\rangle\rightarrow\tilde{\Phi}^{\nu}{}_{\mu_{2},...,\mu_{s},\nu}(x),\qquad\,\,\,\,\,\,M_{12}^{+}|\Phi\rangle\rightarrow g_{\nu(\mu_{1}}\Phi_{\mu_{2},...,\mu_{s})}(x)$